what is the relation between wavelength, frequency, refractive index?

Grade:11

2 Answers

Umakant biswal
7 years ago
v = f. Combining the above expression for velocity with the definition of index of refraction, we find a relationship between the wavelength = v/f in a medium and the wavelength 0 = c/f in vacuum: In the above equation, the frequencies cancel because frequency does not change as light moves from one medium to another.
and the relation between , wavelength and freq is 
freq = c * lambda 
where c is the speed of light 

 
jyoti bhatia
7 years ago
Frequency of light does not change when the light crosses a boundary, only the wavelength changes. Hence, n = c/v = c/(fλ) and thus frequency is inversely proportional to the wavelength.